# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 5.9 Kloosterman’s work
In 2023 R. Kloostermann sent the author the preprint [Klo23] in which among many other things proves that 5.1 is not true for $n\geq 10$. It turns out that for a generic cubic hypersurface $X_{0}$ of dimension $n\geq 10$, we have $T_{0}V_{Z}=T_{0}V_{[Z]}$, even though for Fermat variety this is not true. Actually according to 5.3 at Fermat variety for $n=10$ and $12$ we have the equalities of fourth order and third order neighborhoods, respectively. The equality $T_{0}V_{Z}=T_{0}V_{[Z]}$ can be easily checked by computer using Villaflor’s elegant formula in [VL22b]. I could have done this in 2019 when Villaflor defended his thesis. However, I was too exhausted by my computer calculations in [Mov21a]. The computer code for this verification which uses an example of hypersurface of dimension $10$ in [Klo23] can be obtained in the tex file of the present text in arxiv or the author’s webpage. Recently in [BKU22, Corollary 1.6] the authors have proved that the Hodge loci corresponding to all tensor product of $H^{n}(X_{t},\mathbb{Z})$ and its dual and of positive dimension in the Griffiths period domain is algebraic for $d=3$ and $n\geq 10$. This Hodge loci is formulated in terms of Mumford-Tate groups and it is much larger than the Hodge loci considered in the present text. Even though 5.1 fails for $n\geq 10$, there might be infinite number of special components $V_{i},\ i\in\mathbb{N}$ of Hodge loci in our context which might lie inside a proper algebraic subset of $V$ of ${\sf T}$ (in their terminology maximal element for inclusion). It might be helpful to construct such a $V$ explicitly. If we consider the variation of Hodge structures over $V$ then [BKU22, Corollary 1.6] imply that the generic period domain for $V$ becomes a product of monodromy invariant factors $D_{j}$ and either the level of Hodge decomposition attached to one of these factors is $\leq 2$ or for all except a finite $i\in\mathbb{N}$, the Hodge locus $V_{i}$ has zero dimension projected to one of $D_{j}$. A similar discussion must be also valid for 5.1, that is $n=6,8$, as the Hodge locus $V_{[Z]}$ is atypical in the sense of [BKU22]. For the summary of results on Hodge loci in this general framework see [Kli21].
@@
ˇ$\check{\mathbb{P}}^{\frac{n}{2}}:=\{x_{0}-\zeta_{2d}^{\alpha}x_{1}=x_{2}-\zeta_{2d}^{\alpha}x_{3}=x_{4}-\zeta_{2d}^{\alpha}x_{5}=0\}\cap\mathbb{P}^{\frac{n}{2}+3},$
@@
For $Z$ as in (54), let $V_{[Z]}$, $V_{[\mathbb{P}^{\frac{n}{2}}]}$ and $V_{[\check{\mathbb{P}}^{\frac{n}{2}}]}$ be their corresponding Hodge loci.
@@
In fact, by [Mov21a, Proposition 17.9] we have ${\rm dim}V_{Z}={\rm dim}T_{0}V_{[\mathbb{P}^{\frac{n}{2}}]}\cap T_{0}V_{[\check{\mathbb{P}}^{\frac{n}{2}}]}$ and so we are reduced to show that
@@
By [VL22b, Corollaries 8.2 and 8.3] this is equivalent to show that
@@
(64) $\frac{(I:R_{1}+R_{2})_{3}}{(I:R_{1})_{3}\cap(I:R_{2})_{3}}\simeq\frac{(J^{F}:P_{1}+P_{2})_{3}}{(J^{F}:P_{1})_{3}\cap(J^{F}:P_{2})_{3}}.$
@@
since both are Artin Gorenstein ideals of socle in degree $\frac{n}{2}+4$ (here we use Macaulay theorem [VL22b, Theorem 2.1]) and the right hand side is clearly contained in $(J^{F}:Q)$. In order to prove (64), let $r\in(I:R_{1}+R_{2})_{3}$ such that $r\in(J^{F}:P_{i})_{3}=((J^{F}:Q):R_{i})_{3}$ for both $i=1,2$, then $r\cdot R_{i}\in(J^{F}:Q)\cap\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]=I$ and so $r\in(I:R_{i})_{3}$ for each $i=1,2$. Conversely, given $q\in(J^{F}:P_{1}+P_{2})_{3}$ write it as $q=s+t+u$, where $s\in\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]$, $t\in\langle x_{6}-\zeta_{6}x_{7},x_{8}-\zeta_{6}x_{9},\ldots,x_{n}-\zeta_{6}x_{n+1}\rangle\subseteq\mathbb{C}[x_{0},x_{1},\ldots,x_{n+1}]$ and $u\in\langle x_{7},x_{9},\ldots,x_{n+1}\rangle\subseteq\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\otimes\mathbb{C}[x_{7},x_{9},x_{11},\ldots,x_{n+1}]$. Since $q\cdot(R_{1}+R_{2})\in(J^{F}:Q)$, letting $x_{6}=x_{7}=\cdots=x_{n+1}=0$ it follows that $s\cdot(R_{1}+R_{2})\in I$, i.e. $s\in(I:R_{1}+R_{2})$. On the other hand is clear that $t\in(J^{F}:P_{1})\cap(J^{F}:P_{2})$, then in order to finish the claim it is enough to show that $u\in(J^{F}:P_{1})\cap(J^{F}:P_{2})$. Note that this is clearly true for all monomials appearing in the expansion of $u$ divisible by some $x_{i}^{2}$ for $i>6$ odd. Hence we may assume that
@@
$+\sum_{k>j>i>6\text{ all odd}}p_{ijk}(x_{0},\ldots,x_{5})\cdot x_{i}x_{j}x_{k}.$
@@
$(J^{F}:Q)\cap\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\otimes\mathbb{C}[x_{7},x_{9},x_{11},\ldots,x_{n+1}]=\langle x_{0}^{2},x_{1}^{2},\ldots,x_{5}^{2},x_{7}^{2},x_{9}^{2},\ldots,x_{n+1}^{2}\rangle$
is a monomial ideal. From here it is clear that $u\cdot(R_{1}+R_{2})\in(J^{F}:Q)$ if and only if $p_{i}\cdot(R_{1}+R_{2})\in I$, $p_{ij}\cdot(R_{1}+R_{2})\in I$ and $p_{ijk}\cdot(R_{1}+R_{2})\in I$ for all $k>j>i>6$ odd numbers. Then $p_{i}\in(I:R_{1}+R_{2})_{2}$, $p_{ij}\in(I:R_{1}+R_{2})_{1}$ and $p_{ijk}\in(I:R_{1}+R_{2})_{0}=0$. By [VL22b, Proposition 2.1] we know $(I:R_{1}+R_{2})_{e}=(I:R_{1})_{e}\cap(I:R_{2})_{e}$ for all $e\neq 3$, then $p_{i}\in(I:R_{1})_{2}\cap(I:R_{2})_{2}$ and $p_{ij}\in(I:R_{1})_{1}\cap(I:R_{2})_{1}$ for all $j>i>6$ both odd and so $u\in(J^{F}:P_{1})\cap(J^{F}:P_{2})$ as claimed. This proves (64). Finally, since $(I:R_{1}+R_{2})$, $(I:R_{1})$ and $(I:R_{2})$ are all Artin Gorenstein ideals of socle in degree 3 but they are not equal, we get that $(I:R_{1}+R_{2})_{3}$ is a hyperplane of $\mathbb{C}[x_{0},\ldots,x_{5}]_{3}$ while $(I:R_{1})_{3}\cap(I:R_{2})_{3}$ is a codimension 2 linear subspace of $\mathbb{C}[x_{0},\ldots,x_{5}]_{3}$, hence

### v2

@@ 5.9 Kloosterman’s work
In 2023 R. Kloostermann sent the author the preprint [Kloosterman2023] in which among many other things proves that 5.1 is not true for $n\geq 10$. It turns out that for a generic cubic hypersurface $X_{0}$ of dimension $n\geq 10$, we have $T_{0}V_{Z}=T_{0}V_{[Z]}$, even though for Fermat variety this is not true. Actually according to 5.3 at Fermat variety for $n=10$ and $12$ we have the equalities of fourth order and third order neighborhoods, respectively. The equality $T_{0}V_{Z}=T_{0}V_{[Z]}$ can be easily checked by computer using Villaflor’s elegant formula in [RobertoThesis]. I could have done this in 2019 when Villaflor defended his thesis. However, I was too exhausted by my computer calculations in [ho13]. The computer code for this verification which uses an example of hypersurface of dimension $10$ in [Kloosterman2023] can be obtained in the tex file of the present text in arxiv or the author’s webpage. Recently in [BKU, Corollary 1.6] the authors have proved that the Hodge loci corresponding to all tensor product of $H^{n}(X_{t},\mathbb{Z})$ and its dual and of positive dimension in the Griffiths period domain is algebraic for $d=3$ and $n\geq 10$. This Hodge loci is formulated in terms of Mumford-Tate groups and it is much larger than the Hodge loci considered in the present text. Even though 5.1 fails for $n\geq 10$, there might be infinite number of special components $V_{i},\ i\in\mathbb{N}$ of Hodge loci in our context which might lie inside a proper algebraic subset of $V$ of $\sf T$} (in their terminology maximal element for inclusion). It might be helpful to construct such a $V$ explicitly. If we consider the variation of Hodge structures over $V$ then [BKU, Corollary 1.6] imply that the generic period domain for $V$ becomes a product of monodromy invariant factors $D_{j}$ and either the level of Hodge decomposition attached to one of these factors is $\leq 2$ or for all except a finite $i\in\mathbb{N}$, the Hodge locus $V_{i}$ has zero dimension projected to one of $D_{j}$. A similar discussion must be also valid for 5.1, that is $n=6,8$, as the Hodge locus $V_{[Z]}$ is atypical in the sense of [BKU]. For the summary of results on Hodge loci in this general framework see [Klingler].
@@
$\check{\mathbb{P}}^{\frac{n}{2}}:=\{x_{0}-\zeta_{2d}^{\alpha}x_{1}=x_{2}-\zeta_{2d}^{\alpha}x_{3}=x_{4}-\zeta_{2d}^{\alpha}x_{5}=0\}\cap\mathbb{P}^{\frac{n}{2}+3},$
@@
For $Z$ as in (5.1), let $V_{[Z]}$, $V_{[\mathbb{P}^{\frac{n}{2}}]}$ and $V_{[\check{\mathbb{P}}^{\frac{n}{2}}]}$ be their corresponding Hodge loci.
@@
In fact, by [ho13, Proposition 17.9] we have ${\rm dim}V_{Z}={\rm dim}T_{0}V_{[\mathbb{P}^{\frac{n}{2}}]}\cap T_{0}V_{[\check{\mathbb{P}}^{\frac{n}{2}}]}$ and so we are reduced to show that
@@
By [RobertoThesis, Corollaries 8.2 and 8.3] this is equivalent to show that
@@
(5.11) $\frac{(I:R_{1}+R_{2})_{3}}{(I:R_{1})_{3}\cap(I:R_{2})_{3}}\simeq\frac{(J^{F}:P_{1}+P_{2})_{3}}{(J^{F}:P_{1})_{3}\cap(J^{F}:P_{2})_{3}}.$
@@
since both are Artin Gorenstein ideals of socle in degree $\frac{n}{2}+4$ (here we use Macaulay theorem [RobertoThesis, Theorem 2.1]) and the right hand side is clearly contained in $(J^{F}:Q)$. In order to prove (5.11), let $r\in(I:R_{1}+R_{2})_{3}$ such that $r\in(J^{F}:P_{i})_{3}=((J^{F}:Q):R_{i})_{3}$ for both $i=1,2$, then $r\cdot R_{i}\in(J^{F}:Q)\cap\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]=I$ and so $r\in(I:R_{i})_{3}$ for each $i=1,2$. Conversely, given $q\in(J^{F}:P_{1}+P_{2})_{3}$ write it as $q=s+t+u$, where $s\in\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]$, $t\in\langle x_{6}-\zeta_{6}x_{7},x_{8}-\zeta_{6}x_{9},\ldots,x_{n}-\zeta_{6}x_{n+1}\rangle\subseteq\mathbb{C}[x_{0},x_{1},\ldots,x_{n+1}]$ and $u\in\langle x_{7},x_{9},\ldots,x_{n+1}\rangle\subseteq\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\otimes\mathbb{C}[x_{7},x_{9},x_{11},\ldots,x_{n+1}]$. Since $q\cdot(R_{1}+R_{2})\in(J^{F}:Q)$, letting $x_{6}=x_{7}=\cdots=x_{n+1}=0$ it follows that $s\cdot(R_{1}+R_{2})\in I$, i.e. $s\in(I:R_{1}+R_{2})$. On the other hand is clear that $t\in(J^{F}:P_{1})\cap(J^{F}:P_{2})$, then in order to finish the claim it is enough to show that $u\in(J^{F}:P_{1})\cap(J^{F}:P_{2})$. Note that this is clearly true for all monomials appearing in the expansion of $u$ divisible by some $x_{i}^{2}$ for $i>6$ odd. Hence we may assume that
@@
$\sum_{k>j>i>6\text{ all odd}}p_{ijk}(x_{0},\ldots,x_{5})\cdot x_{i}x_{j}x_{k}.$
@@
$(J^{F}:Q)\cap\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\otimes\mathbb{C}[x_{7},x_{9},x_{11},\ldots,x_{n+1}]=$
$\langle x_{0}^{2},x_{1}^{2},\ldots,x_{5}^{2},x_{7}^{2},x_{9}^{2},\ldots,x_{n+1}^{2}\rangle$
is a monomial ideal. From here it is clear that $u\cdot(R_{1}+R_{2})\in(J^{F}:Q)$ if and only if $p_{i}\cdot(R_{1}+R_{2})\in I$, $p_{ij}\cdot(R_{1}+R_{2})\in I$ and $p_{ijk}\cdot(R_{1}+R_{2})\in I$ for all $k>j>i>6$ odd numbers. Then $p_{i}\in(I:R_{1}+R_{2})_{2}$, $p_{ij}\in(I:R_{1}+R_{2})_{1}$ and $p_{ijk}\in(I:R_{1}+R_{2})_{0}=0$. By [RobertoThesis, Proposition 2.1] we know $(I:R_{1}+R_{2})_{e}=(I:R_{1})_{e}\cap(I:R_{2})_{e}$ for all $e\neq 3$, then $p_{i}\in(I:R_{1})_{2}\cap(I:R_{2})_{2}$ and $p_{ij}\in(I:R_{1})_{1}\cap(I:R_{2})_{1}$ for all $j>i>6$ both odd and so $u\in(J^{F}:P_{1})\cap(J^{F}:P_{2})$ as claimed. This proves (5.11). Finally, since $(I:R_{1}+R_{2})$, $(I:R_{1})$ and $(I:R_{2})$ are all Artin Gorenstein ideals of socle in degree 3 but they are not equal, we get that $(I:R_{1}+R_{2})_{3}$ is a hyperplane of $\mathbb{C}[x_{0},\ldots,x_{5}]_{3}$ while $(I:R_{1})_{3}\cap(I:R_{2})_{3}$ is a codimension 2 linear subspace of $\mathbb{C}[x_{0},\ldots,x_{5}]_{3}$, hence